Find out if the following functions are invertible or not, If it is invertible, then find the rule of the inverse (f^(-1) (x))

1.
f:k → k^+
f(x)=x^2

2.
k^+ → k^+
f(x)=1/x

3.
f:k^+ → k^+
f(x)=x^2
1
Expert's answer
2020-05-25T21:01:54-0400

A function is invertible if the function is surjective and injective.

(More infomartion: https://en.wikipedia.org/wiki/Inverse_function)


Definition. Let "f"  be a function whose domain is a set "X" . The function "f"  is said to be injective provided that for all "a"  and "b"  in "X" , whenever "f(a)=f(b)" , then "a=b" ; that is, "f(a)=f(b)"  implies "a=b" . Equivalently, if "a\neq b" , then "f(a)\neq f(b)" .

Symbolically,



"\displaystyle \forall a,b\in X,\;\;f(a)=f(b)\Rightarrow a=b"

which is logically equivalent to the contrapositive,



"\displaystyle \forall a,b\in X,\;\;f(a)\neq f(b)\Rightarrow a\neq b"

(More infomation: https://en.wikipedia.org/wiki/Injective_function)


Definition. A surjective function is a function whose image is equal to its codomain. Equivalently, a function "f"  with domain "X"  and codomain "Y"  is surjective, if for every "y"  in "Y" , there exists at least one "x"  in "X"  with "\displaystyle f(x)=y" .

Symbolically,

If "\displaystyle f\colon X\rightarrow Y" , then "\displaystyle f"  is said to be surjective if



"\displaystyle \forall y\in Y,\,\exists x\in X,\;\;f(x)=y"

(More information: https://en.wikipedia.org/wiki/Surjective_function)


Hint: For all these questions, I will consider that "k\equiv\reals" .

In our case,

(1) "f : \reals\rightarrow\reals^+;\quad f(x)=x^2"

The function "f(x)=x^2" is not injective in the domain "D(f)=\reals" , since



"(-3)^2=9=3^2,\quad\text{but}\quad -3\neq3"

Conclusion.



"\boxed{f : \reals\rightarrow\reals^+;\quad f(x)=x^2-\text{no inverse function}}"

(2) "f : \reals^+\rightarrow\reals^+;\quad f(x)=1/x"

This function is surjective and injective, therefore the function has an inverse function.



"y=\frac{1}{x}\longrightarrow x=\frac{1}{y}\longrightarrow\\[0.3cm]
\boxed{f(x)=\frac{1}{x}\longrightarrow f^{-1}(x)=\frac{1}{x}}"

Verification,



"f\left(f^{-1}(x)\right)=\frac{1}{\displaystyle\frac{1}{x}}=x"

(3) "f : \reals^+\rightarrow\reals^+;\quad f(x)=x^2"

This function is surjective and injective, therefore the function has an inverse function.



"y=x^2\longrightarrow x=\sqrt{y}\longrightarrow\\[0.3cm]
\boxed{f(x)=x^2\longrightarrow f^{-1}(x)=\sqrt{x}}"

Verification,



"f\left(f^{-1}(x)\right)=\left(\sqrt{x}\right)^2=x"

ANSWER


(1) "f : \reals\rightarrow\reals^+;\quad f(x)=x^2"



"f : \reals\rightarrow\reals^+;\quad f(x)=x^2-\text{no inverse function}"

(2) "f : \reals^+\rightarrow\reals^+;\quad f(x)=1/x"



"f(x)=\frac{1}{x}\longrightarrow f^{-1}(x)=\frac{1}{x}"

(3) "f : \reals^+\rightarrow\reals^+;\quad f(x)=x^2"



"f(x)=x^2\longrightarrow f^{-1}(x)=\sqrt{x}"